17) Find the price p that maximizes revenue if the demand is given by the function q(p) = 500 - 5p
Answer #1

10 = 500-5P RP) = pq@ – 500P-5p² RP) = 500 - lop R80 to find maging ..) 500-80p co - p 50 R(56) = PSODO -1R500 = 12500 9(50)pls like